OpenStudy (anonymous):

Solve for W: P=2L+2W.

OpenStudy (kc_kennylau):

Take away 2L from both sides

jigglypuff314 (jigglypuff314):

you would do that because you want to get "W" all alone "2L" is crowding the "2W"

OpenStudy (anonymous):

2W=P-2L?

jigglypuff314 (jigglypuff314):

yes :) now the "2" in "2W" is crowding the "W" that we want so how would we get rid of that 2 ? :)

OpenStudy (anonymous):

Divide! :)

jigglypuff314 (jigglypuff314):

yes! :D

OpenStudy (anonymous):

W=P-2L/2

OpenStudy (kc_kennylau):

bingo :D

OpenStudy (anonymous):

yaaaaaay!

jigglypuff314 (jigglypuff314):

careful tho, (P-2L)/2 would be correct you need parenthesis

OpenStudy (anonymous):

Thank you...If I simplify it, would it be W=1/2P-L, or I don't need to do that?

jigglypuff314 (jigglypuff314):

you can do that if you would like :) either one would be correct

OpenStudy (anonymous):

Ok thank you both ! Next question!